jason78687:

Car A and B start from rest at the same place at the same time and they travel in same direction. Car A accelerates uniformly at 1m/s^2 upto speed of 36km/h and Car B accelerates at 0.5m/s^2 upto speed of 54km/h. Calculate time and distance at which Car B overtakes A.

Vocaloid:

Use the kinematics equation vf = v0 + at to solve for the time it takes for each car to reach its constant speed (be careful about units since you have both km/hr and m/s units). From there, use delta x = (1/2)at^2 + v0*t to calculate how far each car travels during that time. From there, you can write a distance = (speed)(time) + (the distance travelled so far) for each car, set them equal to each other, and solve for time. Plug time back in to get the distance.

jason78687:

@vocaloid wrote:
Use the kinematics equation vf = v0 + at to solve for the time it takes for each car to reach its constant speed (be careful about units since you have both km/hr and m/s units). From there, use delta x = (1/2)at^2 + v0*t to calculate how far each car travels during that time. From there, you can write a distance = (speed)(time) + (the distance travelled so far) for each car, set them equal to each other, and solve for time. Plug time back in to get the distance.
already tried it doesnt get any better once you start solving

Vocaloid:

Ah I see, I ran the numbers just now and I see what the issue is. I believe car B overtakes A *during* the acceleration period not after. I’ll think a bit more about the approach.

jason78687:

@vocaloid wrote:
Ah I see, I ran the numbers just now and I see what the issue is. I believe car B overtakes A *during* the acceleration period not after. I’ll think a bit more about the approach.
forget it i did it now but it was very complicated